# Evaluating Large Language Model based Personal Information Extraction and Countermeasures

### v1

@@ Abstract
Automatically extracting personal information–such as name, phone number, and email address–from publicly available profiles at a large scale is a stepstone to many other security attacks including spear phishing. Traditional methods–such as regular expression, keyword search, and entity detection–achieve limited success at such personal information extraction. In this work, we perform a systematic measurement study to benchmark large language model (LLM) based personal information extraction and countermeasures. Towards this goal, we present a framework for LLM-based extraction attacks; collect three datasets including a synthetic dataset generated by GPT-4 and two real-world datasets with manually labeled 8 categories of personal information; introduce a novel mitigation strategy based on prompt injection; and systematically benchmark LLM-based attacks and countermeasures using 10 LLMs and our 3 datasets. Our key findings include: LLM can be misused by attackers to accurately extract various personal information from personal profiles; LLM outperforms conventional methods at such extraction; and prompt injection can mitigate such risk to a large extent and outperforms conventional countermeasures. Our code and data are available at: https://github.com/liu00222/LLM-Based-Personal-Profile-Extraction.
Evaluating Large Language Model based Personal Information Extraction and Countermeasures
Yupei Liu Yuqi Jia
Penn State University Duke University
yzl6415@psu.edu yuqi.jia@duke.edu
Jinyuan Jia Neil Zhenqiang Gong
Penn State University Duke University
jinyuan@psu.edu neil.gong@duke.edu
@@ 1 Introduction
Personal information extraction (PIE) refers to automatically extracting personal information, e.g., email address, mailing address, phone number, name, work experience, education background, affiliation, and occupation, from publicly available personal profiles such as personal websites. The extracted personal information can be further leveraged to facilitate other security attacks such as spear phishing [40, 25, 17] and telecommunication fraud [27, 49] at large scale. Traditional PIE methods include regular expression, keyword search, and entity detection [1, 14]. However, these methods achieve limited success at PIE due to insufficient understanding of the semantics in a personal profile, as shown in our experiments.
LLMs, such as GPT-4 [34], PaLM 2 [15], and Gemini [46], represent a significant advancement in natural language processing. Due to their remarkable capability in understanding and generating text, these LLMs have been widely deployed and become conveniently accessible to the general public. For instance, OpenAI provides APIs of ChatGPT to users; and Google provides APIs of PaLM 2 and Gemini to users. These publicly available APIs enable users with limited computational resources to leverage the power of LLMs without the need for extensive infrastructure or technical expertise. This democratization of access has led to a surge in innovative applications, ranging from educational tools and business automation to creative writing and complex problem-solving tasks.
Our work:  In this work, we study how LLMs can be misused by attackers. In particular, we perform the first systematic measurement study to benchmark and quantify the effectiveness of LLM-based PIE attacks (shown in Figure 1) and countermeasures. Towards this goal, our study consists of the following components: formulating an attack framework, collecting datasets, evaluating attacks, formulating countermeasures, and evaluating countermeasures, which we introduce as follows.
@@
Collecting datasets. As LLM-based PIE is an emerging topic, the literature lacks benchmark datasets with ground-truth personal information. To address this challenge, we collect one synthetic dataset by prompting GPT-4 and two real-world datasets by scraping public websites about celebrity and physicians. These datasets contain diverse sets of personal profiles with 8 categories of personal information, such as email address, phone number, mailing address, name, work experience, and education background, that we manually labeled.
Evaluating attacks. We systematically evaluate different LLM-based attacks by configuring the two components of the attack framework using 10 LLMs and our 3 datasets. We have the following key observations. First, LLM-based extraction is effective. For example, GPT-4 achieves 100% and 98% accuracy in extracting email address and phone number in the synthetic dataset. Second, larger LLMs are more successful. For instance, when extracting email address and phone number from the synthetic dataset, vicuna-7b-v1.3 achieves accuracy of 55% and 89% respectively, which are 45% and 9% lower than that of GPT-4. Third, we find that LLMs are more successful at extracting email addresses, phone numbers, mailing addresses, and names while less successful at extracting work experience, educational experience, affiliation, and occupation. Fourth, LLM-based extraction outperforms traditional tools–such as regular expression, keyword search, and entity detection–in almost all scenarios, especially when countermeasures are deployed.
Formulating countermeasures. We first summarize commonly used existing defenses against PIE, including symbol replacement, name replacement, hyperlink, and text-to-image. Moreover, since LLM is vulnerable to prompt injection [23, 29, 42], we propose to leverage prompt injection as a defense to deceive LLMs to extract incorrect personal information. In particular, we can inject a text (serving as injected prompt) into a personal profile; and the injected text is invisible to normal users when the personal profile is rendered as a webpage but will deceive an LLM to output incorrect results when it extracts information from the personal profile. Unlike some existing defenses, e.g., symbol replacement, which can only be applied to a certain category of personal information such as email address, prompt injection is applicable to all categories of personal information. As prompt injection is often viewed as an offensive technique to compromise the security of LLMs, we are the first to show that it can also be used as a defensive technique to protect personal information from being automatically extracted by LLMs.
@@
As a summary, we make the following contributions in this work:
@@
  • We collect 3 datasets with 8 categories of ground-truth personal information.
@@
  • We systematically evaluate various LLM-based PIE attacks and countermeasures using 10 LLMs and our 3 datasets.
Figure 1: LLM-based personal information extraction.
@@ 2.1 Personal Information Extraction (PIE)
Entity detection:  This technique involves identifying and classifying named entities in text. It can often be used for extracting names of organizations, locations, and dates. For instance, entity detection algorithms [1, 14] can identify and classify a university name as an educational institution or a company name as an affiliation.
ML methods:  These methods (e.g., mlscraper [2]) encompass a range of techniques, from simple text parsing to more complex linguistic analysis using machine learning models. They can be used to understand the structure of the personal profiles and the context in which personal information appears.
These traditional methods have two limitations. Firstly, they achieve limited success at extracting personal information especially when extracting certain categories of personal information such as occupation and affiliation. Secondly, they require an attacker to have enough technical background knowledge. For instance, regular expression requires an attacker to have a deep understanding of the regular expression mechanism; and ML methods and entity detection require an attacker to have sufficient understanding of NLP techniques. In contrary, an attacker can simply use off-the-shelf LLMs to accurately extract personal information.
@@ 2.2 Large Language Models (LLMs)
LLMs are a class of neural networks, aiming to process and generate text. In particular, an LLM receives a text, which is commonly referred to as a prompt, and a document (e.g., personal profile) as an input, and generates an output, namely a response. This operation can be expressed as $\hat{y}=f(x,d)$, where $f$ denotes an LLM, $x$ represents the input prompt, $d$ is a document to be analyzed, and $\hat{y}$ represents the generated response. Essentially, the LLM combines $x$ and $d$ as an input and then generates a response. For instance, the prompt could be “Extract email address from the document” and the generated response could be “123@gmail.com”, where the document can be a personal profile from which the attacker aims to extract information. Examples of LLMs include PaLM 2 [15], LLaMA [48], GPT-4 [34], Gemini [46], and many others [21, 47]. Those LLMs are deployed as APIs or open-sourced by model providers. The document can be in different formats, e.g., HTML source code, PDF, markdown, or word documents.
@@ 2.3 Misuse of LLMs
While LLMs offer groundbreaking possibilities in many domains, recent studies revealed that LLMs can be misused by attackers for malicious purposes. First, an attacker can leverage LLMs to generate malicious code or malware [19, 16, 33, 39]. For instance, Carlini [19] showed that, with carefully designed prompts, GPT-4 can be leveraged to automatically break state-of-the-art defenses against adversarial examples [57]. Second, attackers can utilize LLMs to aid the dissemination of misinformation [13, 36, 20, 55]. By generating fake news, articles, or posts on social media, LLMs can be used by attackers to spread false information, which can essentially influence public opinion or even impact the financial market [35]. Third, attackers can misuse LLMs to facilitate the process of writing spear phishing messages [17, 40, 25]. Our work shows another potential misuse of LLMs, in particular for PIE.
@@ 3 Threat Model
Attacker’s goal:  The primary goal of the attacker is to extract personal information from publicly available personal profiles. This includes, but is not limited to, data available on social media platforms, personal homepages, professional networking sites, personal blogs, and websites hosting resumes. The extracted personal information could then facilitate a range of malicious activities. For instance, with the extracted information, attackers can then leverage LLMs to scale spear phishing attacks by automating the process of generating convincing and personalized fake messages to the victims. Another example is spamming call, which targets individuals based on their extracted personal preferences and experience. Additionally, the attacker could undertake the password cracking attempts by leveraging extracted personal information to guess or reset victims’ passwords or security question answers, which could further compromise the privacy and security of individuals.
Attacker’s background knowledge and capabilities:  We consider that an attacker can retrieve publicly available personal profiles, which could include resumes, personal websites, and social media profiles. These profiles can be presented in various data formats such as HTML source code, PDF files, or Word documents. For instance, an attacker can write a crawler to automatically collect such personal profiles on the Internet. Moreover, the attacker has access to an LLM, which could be either an open-source LLM or an API of a closed-source LLM. For instance, with a subscription to the services provided by OpenAI, an attacker can query the GPT-family models (e.g., GPT-3.5, GPT-4) using APIs offered by OpenAI.
@@ 4.1 Defining LLM-based PIE
Attacker-chosen categories of personal information:  We assume that the attacker can arbitrarily select $m$ categories of personal information that the attacker is interested in. For instance, the attacker could select phone number, email address, name, affiliation, and occupation. Given a personal profile $d$, we denote the ground-truth of its $i^{\textit{th}}$ personal information as $y_{i}$, where $i=1,\ldots,m$.
@@ 4.2 Prompt Design
The first key component of the attack framework is the prompt $x$, which is pivotal for LLM-based PIE. The prompt serves as a directive principle to guide the LLM to process and extract specific data from a given personal profile. It plays a crucial role in guiding the LLM to extract personal information accurately and effectively in a format that aligns with the attacker’s expectations. We mainly consider three strategies to construct the prompt: 1) prompt style, 2) in-context learning, and 3) additional instruction to bypass potential defenses in the personal profile.
@@
Table 1: Dataset summary.
Dataset Source Number of profiles Data format Data collection
Synthetic GPT-4 generated 100 HTML Generated by prompting GPT-4
Celebrity List of top-100 famous people [11] 100 HTML Scraped from the Internet
Physician Wikipedia (19th Century American Physicians) [10] 100 HTML Scraped from the Internet
@@ 4.3 Personal Profile Processing
Redundant information filtering:  In this approach, an attacker pre-processes the personal profile using traditional pre-processing techniques for data, which might help filter redundant information. Besides, the LLM may not be able to analyze the original personal profile. In particular, the original profile may contain tokens that are more than what the LLM can take as input, or the original profile contains non-textual information (e.g., images) that a single-modal LLM cannot analyze. For instance, when the personal profile comes in the format of HTML source code, it usually contains many irrelevant information such as CSS functions, style definitions, and HTML tags. This redundant information may make the number of tokens in this profile exceed the maximum input limit of an LLM or distract the LLM from accurately extracting information.
5 Collecting Datasets
To measure the effectiveness of LLM-based PIE attacks and countermeasures, datasets that contain ground-truth personal information are needed. However, the literature lacks such benchmark datasets. To address this challenge, we craft a synthetic dataset and collect two real-world datasets with 8 categories of personal information. A summary of our datasets is presented in Table 1. Next, we introduce the process of collecting the datasets.
@@ 5.1 Synthetic Dataset
We craft a synthetic dataset as the first benchmark dataset of our study. In particular, the synthetic dataset consists of 100 personal profiles. Each personal profile and the corresponding personal information (i.e., the name, email address, mailing address, phone number, occupation, affiliation, work experience, and education background) are uniquely generated using GPT-4 and is in the HTML format, mirroring the typical presentation of online personal profiles. To ensure diversity and realism, we designed a set of prompts that simulate various website styles that are commonly found in realistic online personal profiles. These styles include “neat”, “geeky”, “colorful”, “fancy”, and “formal”. The detailed steps to generate a personal profile are shown as follows:
  • Step 1:
    Randomly draw a style $\mathcal{S}$ from a list of styles, i.e., “neat”, “geeky”, “colorful”, “fancy”, and “formal”.
  • Step 2:
    Query GPT-4 with the prompt shown in Figure 5 in Appendix.
  • Step 3:
    From the GPT-4’s response, we manually extract the personal profile in HTML format and the ground-truth labels for each personal information category.
We generate 100 personal profiles in total. An example of a generated HTML-format personal profile after rendering is shown in Figure 7a in Appendix.
This synthetic dataset brings several benefits to our study. Firstly, it provides a controlled environment to test the effectiveness of LLM-based PIE. Specifically, it allows us to systematically vary the complexity and other parameters of profiles, enabling a comprehensive measurement of LLM-based PIE under different experimental settings. For instance, by prompting GPT-4 with different guiding texts, we are able to generate personal profiles in different styles. Secondly, it offers a benchmark for comparing the performance of different LLMs as well as the performance of traditional PIE methods in a standardized context. Unless otherwise mentioned, we use the synthetic dataset.

### v2

# Evaluating LLM-based Personal Information Extraction and Countermeasures

@@ Abstract
Automatically extracting personal information–such as name, phone number, and email address–from publicly available profiles at a large scale is a stepstone to many other security attacks including spear phishing. Traditional methods–such as regular expression, keyword search, and entity detection–achieve limited success at such personal information extraction. In this work, we perform a systematic measurement study to benchmark large language model (LLM) based personal information extraction and countermeasures. Towards this goal, we present a framework for LLM-based extraction attacks; collect four datasets including a synthetic dataset generated by GPT-4 and three real-world datasets with manually labeled eight categories of personal information; introduce a novel mitigation strategy based on prompt injection; and systematically benchmark LLM-based attacks and countermeasures using ten LLMs and five datasets. Our key findings include: LLM can be misused by attackers to accurately extract various personal information from personal profiles; LLM outperforms traditional methods; and prompt injection can defend against strong LLM-based attacks, reducing the attack to less effective traditional ones.
@@ 1 Introduction
Personal information extraction (PIE) refers to automatically extracting personal information, e.g., email address, mailing address, phone number, name, work experience, education background, affiliation, and occupation, from publicly available personal profiles such as personal websites. The extracted personal information can be further leveraged to facilitate other security attacks such as spear phishing [52, 16] and telecommunication fraud [32, 59] at large scale. Recent spear phishing campaigns, such as the widely publicized incidents targeting WhatsApp accounts and leveraging novel phishing tactics, have heightened public awareness and security concerns [20, 30]. Traditional PIE methods include regular expression, keyword search, and entity detection [1, 13]. However, these methods achieve limited success at PIE due to insufficient understanding of the semantics in a personal profile, as shown in our experiments.
LLMs, such as GPT-4 [41], PaLM 2 [14], and Gemini [58], represent a significant advancement in natural language processing. Due to their remarkable capability in understanding and generating text, these LLMs have been widely deployed and become conveniently accessible to general public. For instance, OpenAI provides ChatGPT APIs to users; and Google provides APIs of PaLM 2 and Gemini to users. These publicly available APIs enable users with limited computational resources to leverage the power of LLMs without the need for extensive infrastructure or technical expertise. This democratization of access has led to a surge in innovative applications, ranging from educational tools and business automation to creative writing and complex problem-solving tasks.
Figure 1: LLM-based personal information extraction.
Our work:  In this work, we study how LLMs can be misused by attackers. We perform the first systematic measurement study to benchmark and quantify the effectiveness of LLM-based PIE attacks (shown in Figure 1) and countermeasures. Towards this goal, our study consists of the following components: formulating an attack framework, collecting datasets, evaluating attacks, formulating countermeasures, and evaluating countermeasures, which we introduce as follows.
@@
Collecting datasets. As LLM-based PIE is an emerging topic, the literature lacks benchmark datasets with ground-truth personal information. To address this challenge, we collect one synthetic dataset by prompting GPT-4 and three real-world datasets by scraping public websites of celebrities, physicians, and computer science professors. These datasets contain diverse sets of personal profiles with 8 categories of personal information, i.e., email address, phone number, mailing address, name, work experience, and education background, that we manually labeled.
Evaluating attacks. We systematically evaluate different LLM-based PIE attacks by configuring the two components of the attack framework using 10 LLMs and 5 datasets (4 datasets collected by us and one benchmark text anonymization dataset introduced in [49]). We have the following observations from our experimental results. First, LLM-based extraction is effective, e.g., GPT-4 achieves 100% and 98% accuracy in extracting email address and phone number in the synthetic dataset. Second, larger LLMs are more successful. For instance, when extracting email address and phone number from the synthetic dataset, vicuna-7b-v1.3 achieves accuracy of 65% and 95% respectively, which are 35% and 3% lower than that of GPT-4. Third, LLMs are more successful at extracting email addresses, phone numbers, mailing addresses, and names while less successful at extracting work experience, educational experience, affiliation, and occupation. Fourth, LLM-based extraction outperforms traditional tools–such as regular expression, keyword search, and entity detection–in almost all scenarios.
Formulating countermeasures. We first summarize commonly used existing defenses against PIE, including symbol replacement, name replacement, hyperlink, and text-to-image. Moreover, since LLM is vulnerable to prompt injection [27, 54], we propose to leverage prompt injection as a defense to deceive LLMs to extract incorrect personal information. In particular, we can inject a text (serving as injected prompt) into a personal profile; and the injected text is invisible to normal users when the personal profile is rendered as a webpage but will deceive an LLM to output incorrect results when it extracts information from the personal profile. Unlike some existing defenses, e.g., symbol replacement, which can only be applied to a certain category of personal information such as email address, prompt injection is applicable to all categories of personal information. As prompt injection is often viewed as an offensive technique to compromise the security of LLMs, we are the first to show that it can also be used as a defensive technique to protect personal information from being automatically extracted by LLMs.
@@
In summary, we make following contributions in this work:
@@
  • We collect 4 datasets with 8 categories of ground-truth personal information.
@@
  • We systematically evaluate LLM-based PIE attacks and countermeasures using 10 LLMs and 5 datasets.
@@ 2.1 Personal Information Extraction (PIE)
Named entity recognition (NER):  This technique involves identifying and classifying named entities in text. It can often be used for extracting names of organizations, locations, and dates. For instance, entity detection algorithms [1, 13] can identify and classify a university name as an educational institution or a company name as an affiliation.
mlscraper:  mlscraper [2] encompass a range of techniques, from simple text parsing to more complex linguistic analysis using machine learning models. They can be used to understand the structure of personal profiles and the context in which personal information appears.
The major limitation of traditional methods [4, 2, 1, 13] is that they achieve limited success at extracting personal information especially when extracting certain categories of personal information such as occupation and affiliation, as shown in our results. In contrary, an attacker can use off-the-shelf LLMs to accurately extract personal information.
LLM-based PIE:  Bubeck et al. [18] evaluated LLM-based PIE on short/simple texts for non-security-critical categories of personal information. Sancheti et al. [51] aims to consolidate the web profiles of an individual into a single one. Staab et al. [53] also studied LLM-based PIE attacks. Compared with [18, 51, 53], first, we focus on different types of personal information, and some of them in our work are arguably more security-critical, e.g., email addresses are highly relevant for phishing attacks. Second, we study different ways of using LLM, e.g., prompt styles, for PIE, while [18, 51, 53] do not. Third, we evaluate multiple defenses, which were not covered in [18, 51, 53]. For example, [53] studies anonymization (e.g., replacing email address with ******) as a defense, which substantially reduces utility as it makes personal information unusable. As a comparison, we study more practical and widely deployed defenses, such as symbol and keyword replacement. Furthermore, we demonstrate that prompt injection combined with traditional defenses can defend against existing PIE attacks. We note that some other studies extract personal information from the training data memorized by LLMs [36, 40, 45]. They are not applicable to extract personal information from data that are not in the training data.
Text anonymization:  Text anonymization aims to obfuscate personal information in texts, which also involves personal information detection. Existing studies [49, 29, 43, 33] mainly use traditional PIE techniques, e.g., NER methods based on spaCy/BERT. A recent work [46] explored the use of LLMs, but its method directly uses LLMs to anonymize/deanonymize text without detecting personal information.
@@ 2.2 Large Language Models (LLMs)
LLMs are a class of neural networks, aiming to process and generate text. In particular, an LLM receives a text, which is commonly referred to as a prompt, and a document (e.g., personal profile) as an input, and generates an output, namely a response. This operation can be expressed as $\hat{y}=f(x,d)$, where $f$ denotes an LLM, $x$ is the input prompt, $d$ is a document to be analyzed, and $\hat{y}$ is the generated response. Essentially, the LLM combines $x$ and $d$ as an input and then generates a response. For instance, the prompt could be “Extract email address from the document” and the generated response could be “123@gmail.com”, where the document can be a personal profile from which the attacker aims to extract information. Examples of LLMs include GPT-4 [41], Gemini [58], and many others [25, 14, 23, 6]. Those LLMs are deployed as APIs or open-sourced. The document can be in different formats, e.g., HTML or PDF.
@@ 2.3 Misuse of LLMs
While LLMs offer groundbreaking possibilities in many domains, recent studies revealed that LLMs can be misused by attackers for malicious purposes. First, an attacker can leverage LLMs to generate malicious code or malware [19, 15, 39, 50]. For instance, Carlini [19] showed that, with carefully designed prompts, GPT-4 can be leveraged to automatically break state-of-the-art defenses against adversarial examples [65]. Second, attackers can utilize LLMs to aid the dissemination of misinformation [12, 44, 21, 63]. By generating fake news, articles, or posts on social media, LLMs can be used by attackers to spread false information, which can essentially influence public opinion or even impact the financial market [42]. Third, attackers can misuse LLMs to facilitate the process of writing spear phishing messages [16, 52]. Our work shows another potential misuse of LLMs, in particular for PIE.
@@ 3 Threat Model
Attacker’s goal:  The primary goal of the attacker is to extract personal information at scale from publicly available personal profiles using automated techniques. This includes, but is not limited to, data available on social media platforms, personal homepages, professional networking sites, personal blogs, and websites hosting resumes. The extracted personal information could then facilitate a range of malicious activities. For instance, with the extracted information, attackers can then leverage LLMs to scale spear phishing attacks by automating the process of generating convincing and personalized fake messages to the victims. Another example is spamming call, which targets individuals based on their extracted personal preferences and experience. Additionally, the attacker could undertake the password cracking attempts by leveraging extracted personal information to guess or reset victims’ security question answers, which could further compromise the privacy and security of individuals.
Attacker’s background knowledge and capabilities:  We consider that an attacker can retrieve publicly available personal profiles, which could include resumes, personal websites, and social media profiles. These profiles can be presented in various data formats such as HTML or PDF. For instance, an attacker can write crawlers to automatically collect such profiles on the internet. Given the personal profiles, we consider an attacker can use automated PIE methods to extract personal information. For instance, an attacker can use LLM-based PIE methods, where we consider an attacker has access to an open-source LLM or an API of a closed-source LLM. As an example, with a subscription to the services provided by OpenAI, an attacker can query the GPT models (e.g., GPT-3.5, GPT-4) using OpenAI’s API.
@@ 4.1 Defining LLM-based PIE
Attacker-chosen categories of personal information:  We assume that the attacker can arbitrarily select $m$ categories of personal information that the attacker is interested in. For instance, the attacker could select phone number, email address, name, affiliation, and occupation. Given a personal profile $d$, we denote the ground-truth of its $i^{\textit{th}}$ personal information as $y_{i}$, where $i=1,2,\ldots,m$.
@@ 4.2 Prompt Design
The first key component of the attack framework is the prompt $x$, which is pivotal for LLM-based PIE. The prompt serves as a directive principle to guide the LLM to process and extract specific data from a given personal profile. We mainly consider three strategies to construct the prompt: 1) prompt style, 2) in-context learning, and 3) additional instruction to bypass potential defenses in the personal profile.
@@ 4.3 Personal Profile Processing
Redundant information filtering:  In this approach, an attacker pre-processes the personal profile using traditional pre-processing techniques for data, which might help filter redundant information. Besides, the LLM may not be able to analyze the original personal profile. In particular, the original profile may contain tokens that are more than what the LLM can take as input, or the original profile contains non-textual information (e.g., images) that a single-modal LLM cannot analyze. For instance, when the personal profile comes in the format of HTML source code, it usually contains irrelevant information such as CSS functions, style definitions, and HTML tags, which may make the number of tokens in this profile exceed the maximum input limit of an LLM or distract the LLM from accurately extracting information.
5 Datasets
Table 1: Dataset summary.
Dataset Source #Profiles Data format Data collection
Synthetic GPT-4 generated 100 HTML Generated by prompting GPT-4
Celebrity List of top-100 famous people [10] Scraped from the Internet
Physician Wikipedia (19th Century American Physicians) [9]
Professor Profiles of EECS professors in Berkeley, MIT, Stanford, and CMU
Court English court cases from the European Court of Human Rights 127 Text Existing work [49]
To measure the effectiveness of PIE attacks and defenses, datasets that contain ground-truth personal information are needed. However, the literature lacks such datasets. In response, we craft a synthetic dataset and collect three real-world datasets with 8 categories of personal information.
@@ 5.1 Synthetic Dataset
We craft a synthetic dataset as the first benchmark dataset of our study. In particular, the synthetic dataset consists of 100 personal profiles. Each personal profile and the corresponding personal information (i.e., the name, email address, mailing address, phone number, occupation, affiliation, work experience, and education background) are uniquely generated using GPT-4 and is in the HTML format, mirroring the typical presentation of online personal profiles. To ensure diversity and realism, we designed a set of prompts that simulate various website styles that are commonly found in realistic online personal profiles. These styles include “neat”, “geeky”, “colorful”, “fancy”, and “formal”. The detailed steps to generate a personal profile are as follows: 1) Randomly draw a style $\mathcal{S}$ from a list of styles, i.e., “neat”, “geeky”, “colorful”, “fancy”, and “formal”; 2) Query GPT-4 with the prompt shown in Figure 6 in the Appendix; and 3) From the GPT-4’s response, we manually extract the personal profile in HTML format and the ground-truth labels for each personal information category. We generate 100 personal profiles in total. An example of a generated personal profile is shown in Figure 5 in the Appendix.
This synthetic dataset brings several benefits to our study. Firstly, it provides a controlled environment to test the effectiveness of LLM-based PIE. Specifically, it allows us to systematically vary the complexity and other parameters of profiles, enabling a comprehensive measurement of LLM-based PIE under different experimental settings. For instance, by prompting GPT-4 with different guiding texts, we are able to generate personal profiles in different styles. Secondly, it offers a benchmark for comparing the performance of different LLMs as well as the performance of traditional PIE methods in a standardized context.